# Benchmarking Vision Language Models for Cultural Understanding

### v2

@@ Abstract
Foundation models and vision-language pre-training have notably advanced Vision Language Models (VLMs), enabling multimodal processing of visual and linguistic data. However, their performance has been typically assessed on general scene understanding – recognizing objects, attributes, and actions – rather than cultural comprehension. This study introduces CulturalVQA, a visual question-answering benchmark aimed at assessing VLM’s geo-diverse cultural understanding. We curate a collection of 2,378 image - question pairs with 1-5 answers per question representing cultures from 11 countries across 5 continents. The questions probe understanding of various facets of culture such as clothing, food, drinks, rituals, and traditions. Benchmarking VLMs on CulturalVQA, including GPT-4V and Gemini, reveals disparity in their level of cultural understanding across regions, with strong cultural understanding capabilities for North America while significantly lower performance for Africa. We observe disparity in their performance across cultural facets too, with clothing, rituals, and traditions seeing higher performances than food and drink. These disparities help us identify areas where VLMs lack cultural understanding and demonstrate the potential of CulturalVQA as a comprehensive evaluation set for gauging VLM progress in understanding diverse cultures.
[Uncaptioned image]    https://culturalvqa.org/
@@
Shravan Nayak1,2     Kanishk Jain1,2     Rabiul Awal1,2 Siva Reddy1,3     Sjoerd van Steenkiste4     Lisa Anne Hendricks4 Karolina Stańczak1,3     Aishwarya Agrawal1,2 1Mila – Quebec AI Institute,   2Université de Montréal,   3McGill University,   4Google DeepMind Correspondence: shravan.nayak@mila.quebec
@@ 1 Introduction
Recent multimodal vision-language models (VLMs) (Radford et al., 2021; Liu et al., 2023; Peng et al., 2023; Chen et al., 2024; Lu et al., 2024) have shown impressive performance in tasks such as image-to-text generation, visual question answering, and image captioning, inter alia. These tasks predominantly focus on general scene understanding capabilities such as recognizing attributes, objects, and actions in scenes containing objects in their common context (Lin et al., 2014). However, given the advancing capabilities of VLMs, we believe the time is now ripe to hold VLMs to higher standards. We believe that to support increasingly global digital interactions, VLMs must also be capable of understanding the cultural values (Liu et al., 2021) such as beliefs, rituals, and traditions, for a variety of cultures in the world.
@@
Figure 1: Samples from CulturalVQA (top). The performance of VLMs over time, segmented by non-Western (red) and Western (blue) countries, with model release dates annotated (bottom). Dashed and solid lines differentiate trends for non-Western and Western countries, respectively. VLMs’ understanding of non-Western cultures has been in a steep upward trend since Jan ’24, LLaVa-Next Liu et al. (2024) release.
@@
In order to adequately assess whether the current state-of-the-art VLMs – including proprietary models such as GPT-4V (OpenAI, 2023) and Gemini (Gemini Team et al., 2023) – encode cultural knowledge, we need systematic benchmarks. However, evaluating cultural understanding is a challenging task since culture is a multifaceted concept consisting of both tangible (e.g., clothing, and food) as well as intangible elements (e.g., ritual practices). Current benchmarks in this domain, including MaRVL (Liu et al., 2021) and GD-VCR (Yin et al., 2021), while offering foundational insights, have critical shortcomings. MaRVL primarily focuses on visual reasoning tasks (e.g., counting, spatial reasoning) on top of images sourced from various cultures, and lacks probing cultural common sense – the knowledge bank shared by the members of a cultural group (see § 3). GD-VCR although explores commonsense, it is limited by its reliance on movie scenes, which do not encompass the broader spectrum of everyday cultural contexts.
@@
In response to the above challenges, we propose CulturalVQA, a novel benchmark specifically designed to assess cultural understanding of VLMs. CulturalVQA is based on Visual Question Answering (VQA), requiring models to integrate both visual and textual information, which permits the formulation of diverse questions, thereby enabling the evaluation of a model’s understanding of cultural nuances. The CulturalVQA benchmark extends the language-only CANDLE dataset (Nguyen et al., 2023), which provides a comprehensive collection of cultural commonsense knowledge assertions. We expanded this dataset by automatically collecting images that depict the cultural concept described by the assertions. On top of these images, we collect questions and answers by employing annotators from different cultures who would be familiar with the different cultural concepts depicted in the images. See Fig. 1 (top) for some examples of questions and answers. Our benchmark consists of 2,378 questions collected on top of 2,328 unique images with 1-5 answers per question (total 7,206 answers) from 11 countries. [^1] We provide a data statement in § A. We also present several analyses to better understand the nature of questions and answers in our benchmark.
@@
Further, we systematically evaluate several state-of-the-art VLMs on CulturalVQA. Our evaluation reveals a distinct performance gap between proprietary and open-source models, with open-source models significantly underperforming in comparison (e.g., there is a 29.78% gap between the highest-performing closed-source model and its open-source counterpart in the country for which the models perform the worst, Ethiopia). Additionally, we observe a significant disparity in model performance across countries. For instance, the highest-performing proprietary model, GPT-4V, achieves 67% and 72% accuracy for North American cultural concepts while only between 43% and 56% accuracy on concepts from Africa. VLMs also show varying degrees of proficiency across cultural facets, with closed-source VLMs performing better on questions about rituals and traditions while scoring worse on those related to clothing, food, and drink. We develop CulturalVQA as a comprehensive evaluation set for gauging VLMs’ progress in understanding diverse cultures and highlighting areas where VLMs lack cultural understanding, with the hope that our benchmark will contribute to accelerating the advancements of VLMs in their cultural understanding, as illustrated in Fig. 1.
@@ 2 Related work
Cultural understanding is closely related to geo-diverse understanding. Existing geo-diverse datasets, for instance, the Dollar Street dataset (Gaviria Rojas et al., 2022) includes 38,479 images of everyday household items from homes around the world, while the GLDv2 dataset (Weyand et al., 2020) contains 5 million images and 200k distinct instance labels of natural and human-made landmarks, but both only test recognition capabilities as opposed to cultural understanding. The GD-VCR dataset (Yin et al., 2021) probes cultural understanding, but its reliance on cinematic scenes limits the diversity of real-world cultural contexts it can have. Another related line of work focuses on multilingual understanding. For instance, Bugliarello et al. (2022) bring together five datasets across a number of tasks in 20 languages. However, their focus lies in multilingual understanding (as opposed to cultural understanding). Another multilingual dataset, MaRVL (Liu et al., 2021), tests visually grounded reasoning across multiple languages and cultures. However, MaRVL does not explore the cultural common sense of rituals and traditions. Additionally, the XM3600 dataset (Thapliyal et al., 2022), includes image captions from 36 regions and languages, thus providing a broad geographical coverage but nonetheless contains mostly Western content and lacks depth in the included cultural concepts (Pouget et al., 2024). Closest to our work, the MaXM benchmark (Changpinyo et al., 2023), building on the XM3600 dataset, and the concurrent study by Romero et al. (2024) both use the VQA format to explore regional and cultural understanding. MaXM focuses primarily on the ability to process images from varied regions rather than on nuanced cultural understanding. Romero et al. (2024) study cultural questions in a multilingual setup. However, their focus diverges from ours as, like MaRVL, they allocate a much smaller proportion of their dataset to traditions and rituals.
@@ 3 CulturalVQA: Dataset Creation
Culture is a multifaceted concept that describes the way of life of a collective group of people, distinguishing them from other groups with different cultures (Hofstede et al., 2010; Hershcovich et al., 2022). In this paper, we use the concept of a country as a proxy for a cultural group (Adilazuarda et al., 2024).  [^2] See § 7 for a discussion of this choice. Our work assumes common ground within a cultural group by probing culturally relevant concepts that are collectively understood, as well as shared cultural common sense employed in reasoning (Hershcovich et al., 2022). For instance, lavash – a traditional Persian bread (see Fig. 2) – is an example of a culturally relevant concept, while the common practice of waltzing at weddings exemplifies the cultural common sense among Germans.
@@
The image selection begins with the CANDLE dataset (Nguyen et al., 2023), which provides a rich collection of Cultural Commonsense Knowledge (CCSK). Each of the 1.1 million entries includes URLs to webpages with relevant CCSK data from the C4 corpus (Raffel et al., 2020). Inspired by findings from (Zhu et al., 2023), which highlighted that 80% of webpages in the C4 corpus contain relevant images, we scrape images from these URLs, focusing particularly on CCSK data from the geography and religion domains of our selected countries.
To refine the image dataset derived from web scraping, we applied filters for aspect ratio, size, and specific keywords, and used CLIP similarity (Hessel et al., 2021) to rank images for cultural relevance. Images with low CLIP scores were discarded, and we sampled the remaining images based on their scores, with higher scores having a higher probability of selection. Details of the image filtering process can be found in § B.
Question Collection
Following the conceptual culture framework by Hofstede et al. (2010), we directed annotators to create questions that are easily answerable by someone from their own culture but challenging for outsiders. To elicit such questions, annotators were guided by the instructions shown in § C and were provided with images and additional context to cultural concepts presented in the image (retrieved from CANDLE). We encouraged them to create questions based on their cultural knowledge, using the additional context (accessible behind a click-to-expand box) only when absolutely necessary. Annotators were also advised to skip images if they found them culturally irrelevant or were unfamiliar with the depicted content.
Initially, for this task, we attempted to engage professional annotators from the Amazon Mechanical Turk (MTurk) platform. However, we encountered challenges in finding sufficient presence of annotators from some of the targeted countries. Therefore, we expanded our search to other communities with a broad cultural representation, including Masakhane, an African NLP organization, and Mila, an international academic AI research institute. Employing annotators from these sources, we conducted pilot studies to iterate over the task instructions and to pre-select high-quality participants.
Answer Collection
Next, we asked the annotators to write answers to the questions created in the previous step, ensuring that the answers reflected common agreement within their culture (see instructions in § D). We prompt them to use English for universal concepts like cats or apples and use widely recognized and agreed upon local terms for concepts like beliefs, festivals, or local cuisine, rather than translating these terms into English. For example, the annotators should write the term naan instead of Indian bread. This approach preserves the cultural specificity of the collected answers. Further, we instructed annotators to be as precise as possible in their answers (e.g., sushi instead of food and Oolong tea instead of tea) and to keep their responses concise, ideally between one to three words.
4 Dataset Analysis
This section provides a detailed analysis of our dataset’s composition and characteristics. In particular, we offer an analysis of images, questions, answers, and cultural concepts included in the CulturalVQA dataset.
@@ 4 Dataset Analysis
Our dataset comprises of 2,328 unique images. In Fig. 2, we show representative samples showcasing the images and cultural concepts within our dataset. The concepts depicted in the images are sourced from 11 countries, selected through a strategic process to ensure extensive cultural representation. The distribution of unique image count per country is detailed in Fig. 3.
@@
We collected 2,378 questions in total. In Fig. 3, we present the number of unique questions per country. The questions have an average length of 10.98 words (see Fig. 3 for country-wise breakdown). Most frequent question types include ‘What’(51.3%), ‘Which’(11.2%), ‘In’ (5.6%), ‘Why’ (3.4%),‘Where’(3.1%) ‘Identify’(3.0%), and ‘How’ (2.7%) questions. For example, ‘What’ questions often relate to identifying cultural entities like saree or Dirndl (traditional Indian and German dresses, respectively) in the clothing category, or festivals like Ramadan (observed e.g., in Nigeria) and Spring Festival (celebrated in China) among rituals. ‘Where’ questions inquire about locations significant to specific foods, such as the origins of Quebec chicken. Finally, we analyzed whether the collected questions contain stereotypes and found that they are largely absent (see § E).
@@
CulturalVQA consists of 7,206 manually curated answers in total. [^4] We collected 1-5 answers per question, depending on the availability of annotators. The average answer length is 1.73 words (see Fig. 3 (B) for country wise breakdown). We assess whether answers predominantly feature terms from local languages. To this end, we verified how many answers have corresponding English Wikipedia titles; for 80% of the answers at least one of the answer words is contained in at least one Wikipedia title. Thus our benchmark is still suitable for English VLMs.
@@
According to the pie chart in Fig. 4, food-related questions are most prevalent, accounting for 37.3% of the dataset, followed closely by traditions and rituals, which represent 26.1% and 18% respectively. Thus, roughly 44% of the questions in our dataset probe for cultural understanding of the intangible aspects of culture (rituals and traditions)! The word clouds generated from the collected answers in Fig. 4 reveal diverse expressions of rituals and traditions represented by terms like hamam (Turkey) and meskel (Ethiopia). Further, the food category includes diverse items such as feijoada (Brazil), fufu (Nigeria), and vada (India) indicating a geo-diverse culinary scope. While the clothing category is the least prevalent in the dataset, it shows the highest variety in terms of collected answers. The drink category is notably one of the smallest, both in terms of the size and number of unique answers.
5 Benchmarking VLMs on CulturalVQA

### v3

@@ Abstract
Foundation models and vision-language pre-training have notably advanced Vision Language Models (VLMs), enabling multimodal processing of visual and linguistic data. However, their performance has been typically assessed on general scene understanding – recognizing objects, attributes, and actions – rather than cultural comprehension. This study introduces CulturalVQA, a visual question-answering benchmark aimed at assessing VLM’s geo-diverse cultural understanding. We curate a collection of 2,378 image - question pairs with 1-5 answers per question representing cultures from 11 countries across 5 continents. The questions probe understanding of various facets of culture such as clothing, food, drinks, rituals, and traditions. Benchmarking VLMs on CulturalVQA, including GPT-4o and Gemini, reveals disparity in their level of cultural understanding across regions, with strong cultural understanding capabilities for North America while significantly lower performance for Africa. We observe disparity in their performance across cultural facets too, with clothing, rituals, and traditions seeing higher performances than food and drink. These disparities help us identify areas where VLMs lack cultural understanding and demonstrate the potential of CulturalVQA as a comprehensive evaluation set for gauging VLM progress in understanding diverse cultures.
[Uncaptioned image]    https://culturalvqa.org
@@
Shravan Nayak1,2     Kanishk Jain1,2     Rabiul Awal1,2 Siva Reddy1,3     Sjoerd van Steenkiste4     Lisa Anne Hendricks5 Karolina Stańczak1,3     Aishwarya Agrawal1,2 1Mila – Quebec AI Institute,  2Université de Montréal,  3McGill University, 4Google Research,  5Google DeepMind Correspondence: shravan.nayak@mila.quebec
@@ 1 Introduction
Recent multimodal vision-language models (VLMs) (Radford et al., 2021; Liu et al., 2023; Peng et al., 2023; Chen et al., 2024; Lu et al., 2024) have shown impressive performance in tasks such as image-to-text generation, visual question answering, and image captioning, inter alia. These tasks predominantly focus on general scene understanding capabilities such as recognizing attributes, objects, and actions in scenes containing objects in their common context (Lin et al., 2014). However, given the advancing capabilities of VLMs, we believe the time is now ripe to hold VLMs to higher standards. Indeed, to support increasingly global digital interactions, VLMs must also be capable of understanding the cultural values (Liu et al., 2021) such as beliefs, rituals, and traditions, for a variety of cultures in the world.
@@
Figure 1: The performance of VLMs over time, segmented by non-Western (red) and Western (blue) countries, with model release dates annotated (bottom). Dashed and solid lines differentiate trends for non-Western and Western countries respectively. VLMs’ understanding of non-Western cultures has been in a steep upward trend since Jan ’24.
@@
In order to adequately assess whether the current state-of-the-art VLMs – including proprietary models such as GPT-4o (OpenAI, 2023) and Gemini (Gemini Team et al., 2023) – encode cultural knowledge, we need systematic benchmarks. However, evaluating cultural understanding is a challenging task since culture is a multifaceted concept consisting of both tangible (e.g., clothing, and food) as well as intangible elements (e.g., ritual practices). Current benchmarks in this domain, including MaRVL (Liu et al., 2021) and GD-VCR (Yin et al., 2021), while offering foundational insights, have critical shortcomings. MaRVL primarily focuses on visual reasoning tasks (e.g., counting, spatial reasoning) on top of images sourced from various cultures, and lacks probing cultural common sense – the knowledge base shared by the members of a cultural group (see § 3). While GD-VCR does consider commonsense to a degree, it primarily considers movie scenes, which do not encompass the broader spectrum of everyday cultural contexts.
@@
In response to the above challenges, we propose CulturalVQA, a novel benchmark specifically designed to assess cultural understanding of VLMs. CulturalVQA is based on Visual Question Answering (VQA), requiring models to integrate both visual and textual information, which permits the formulation of diverse questions, thereby enabling the evaluation of a model’s understanding of cultural nuances. The CulturalVQA benchmark extends the language-only CANDLE dataset (Nguyen et al., 2023), which provides a comprehensive collection of cultural commonsense knowledge assertions. We expand this dataset by automatically collecting images that depict the cultural concept described by the assertions. On top of these images, we collect questions and answers by employing annotators from different cultures who would be familiar with the different cultural concepts depicted in the images. See Fig. 2 for some examples of questions and answers. Our benchmark consists of 2,378 questions collected on top of 2,328 unique images with 1-5 answers per question (total 7,206 answers) from 11 countries. [^1] We provide a data statement in § A. We also present several analyses to better understand the nature of questions and answers in our benchmark.
Dataset No. Regions No. Questions No. Images Multilingual? Task Format Culturally Diverse Images? Nature of Questions MaXM (Changpinyo et al., 2023) 7 2142 335 Yes Open-ended No (Pouget et al., 2024) General reasoning GDVCR (Yin et al., 2021) 4 886 328 No Multiple choice Yes (movie scenes only) Cultural understanding MaRVL (Liu et al., 2021) 5 5670 4914 Yes True/False Yes Cultural reasoning CVQA (Romero et al., 2024) 28 9044 4560 Yes Multiple choice Yes Cultural understanding CulturalVQA (Ours) 11 2378 2328 No Open-ended Yes Cultural understanding
Table 1: Comparison of various datasets closely related to CulturalVQA across different axes.
@@
Further, we systematically evaluate several state-of-the-art VLMs on CulturalVQA. Our evaluation reveals a distinct performance gap between proprietary and open-source models, with open-source models significantly underperforming in comparison (e.g., there is a 29.78% gap between the highest-performing closed-source model and its open-source counterpart in the country for which the models perform the worst, Ethiopia). Additionally, we observe a significant disparity in model performance across countries. For instance, the highest-performing proprietary model, GPT-4o, achieves 67% and 72% accuracy on North American cultural concepts while only between 43% and 56% accuracy on concepts from Africa. VLMs also show varying degrees of proficiency across cultural facets, with closed-source VLMs performing better on questions about rituals and traditions while scoring worse on those related to clothing, food, and drink. We develop CulturalVQA as a comprehensive evaluation set for gauging VLMs’ progress in understanding diverse cultures and highlighting areas where VLMs lack cultural understanding. We hope that our benchmark will contribute to accelerating the advancements of VLMs in their cultural understanding, as illustrated in Fig. 1.
@@ 2 Related work
Cultural understanding is closely related to geo-diverse understanding. For instance, the Dollar Street dataset (Gaviria Rojas et al., 2022) includes 38,479 images of everyday household items from homes around the world, while the GLDv2 dataset (Weyand et al., 2020) contains 5 million images and 200k distinct instance labels of natural and human-made landmarks, but both only test recognition capabilities as opposed to cultural understanding. Burda-Lassen et al. (2024) introduce Mosaic-1.5k, a culture-specific captioning dataset that includes images from various regions. Bhatia et al. (2024) propose GLobalRG, which aims to evaluate retrieval and grounding capabilities in VLMs across 15 and 50 countries respectively. Another related line of work focuses on multilingual understanding. For instance, Bugliarello et al. (2022) unify five datasets across a number of tasks in 20 languages. However, their focus lies in multilingual understanding as opposed to cultural understanding. Additionally, the XM3600 dataset (Thapliyal et al., 2022), includes image captions from 36 regions and languages, but lacks depth in cultural concepts, making it insufficient for evaluating cultural diversity in VLMs (Pouget et al., 2024).
Closest to our work are the following benchmarks: MaXM (Changpinyo et al., 2023), GD-VCR (Yin et al., 2021), MaRVL (Liu et al., 2021) and the concurrent work CVQA (Romero et al., 2024). MaXM lacks depth in cultural concepts, as it builds on XM3600 images. Also, its questions focus more on reasoning and general image understanding rather than cultural understanding  [^2] We manually annotated 100 random questions from the English subset of the MaXM and found the following distribution: color - 3.7%, spatial understanding - 12.9%, scene understanding - 42.6%, Yes/No - 20.4%, counting - 20.4% . The GD-VCR dataset probes cultural understanding, but its reliance on cinematic scenes limits the diversity of real-world cultural contexts it can have. Moreover, they rely on a multiple-choice evaluation format, which can be influenced by the difficulty of answer choices. We believe an open-ended evaluation provides a more faithful assessment of the models’ underlying capabilities. Similarly, while MaRVL tests visually grounded reasoning across multiple languages and cultures, it does not assess cultural common sense related to rituals and traditions and also employs a True/False evaluation style. CVQA studies cultural questions in a multilingual setup. However, their focus diverges from ours as they allocate a much smaller proportion of their dataset to traditions and rituals (13% as compared to 44.1 % in CulturalVQA) and use a multiple-choice evalaution format. A comprehensive comparison of different datasets across various dimensions is presented in Tab. 1. CulturalVQA uniquely emphasizes open-ended evaluation, includes culturally diverse images (i.e., images from multiple cultures), and its questions probe for cultural understanding by design. The combination of these characteristics sets CulturalVQA apart from other datasets that either lack culturally diverse images (MaXM), or use restricted evaluation formats such as multiple-choice (CVQA, GDVCR) or True/False (MaRVL).
@@ 3 CulturalVQA: Dataset Creation
Culture is a multifaceted concept that describes the way of life of a collective group of people, distinguishing them from other groups with different cultures (Hofstede et al., 2010; Hershcovich et al., 2022). In this paper, we use the concept of a country as a proxy for a cultural group (Adilazuarda et al., 2024) [^3] See § 7 for a discussion of these choices.. Our work assumes common ground within a cultural group by probing culturally relevant concepts that are collectively understood, as well as shared cultural common sense employed in reasoning (Hershcovich et al., 2022). For instance, lavash – a traditional Persian bread (see Fig. 2) – is an example of a culturally relevant concept, while the common practice of waltzing at weddings exemplifies the cultural common sense among Germans.
@@
We use the CANDLE dataset (Nguyen et al., 2023) for our image source which contains 1.1 million entries of Cultural Commonsense Knowledge (CCSK) along with URLs to corresponding webpages from the C4 corpus (Raffel et al., 2020). The CANDLE dataset represents cultural concepts from approximately 196 countries and 80% of web pages in this corpus contain images related to the text (Zhu et al., 2023). This allows us to begin with a culturally relevant pool of images.
We apply filters for aspect ratio, size, and specific keywords to refine the image dataset. Further, we use CLIP similarity (Hessel et al., 2021) to filter images for cultural relevance, discarding those with a CLIP score below a threshold determined through qualitative evaluation of sample images. [^5] Threshold of 23 (precision = 0.92, recall = 0.96) Since our initial pool already contains culturally relevant images, there is minimal risk of introducing western-centric biases through the use of CLIP, despite potential biases in its pretraining data. To further ensure quality, we apply an additional round of human filtering (detailed in the next section). Thus, our multi-stage filtering ensures that the final set of images is appropriate for cultural annotations. Further details of the image filtering process are provided in § B.
@@
Question Collection
Following the conceptual culture framework by Hofstede et al. (2010), we direct annotators to create questions that are easily answerable by someone from their own culture but challenging for outsiders. To elicit such questions, we provide annotators the instructions shown in § N as well as images and additional context for the cultural concepts present in the image (retrieved from CANDLE). We encourage them to create questions based on their own cultural knowledge, using the additional context (accessible behind a click-to-expand box) only when absolutely necessary. We also advise annotators to skip images if they found them culturally irrelevant or were unfamiliar with the depicted content. This adds an additional layer of filtering, resulting in annotators discarding 19.64% of the images shown to them.
Answer Collection
Next, we ask the annotators to write answers to the questions created in the previous step, while ensuring that the answers reflected common agreement within their culture (see instructions in § O). Here we prompt them to use English for universal concepts like cats or apples and use widely recognized and agreed upon local terms for concepts like festivals or local cuisine, rather than translating them into English. For example, the annotators should write the term Naan instead of Indian bread. This approach preserves the cultural specificity of the collected answers. Further, we instruct annotators to be as precise as possible in their answers (e.g., sushi instead of food and Oolong tea instead of tea) and to keep their responses concise, ideally between one to three words.
Further details about the rationale behind the data curation process and the challenges encountered are provided in § I.
4 Dataset Analysis
This section provides a detailed analysis of CulturalVQAs’ composition and characteristics including analysis of images, questions, answers, and cultural concepts contained in it.
@@ 4 Dataset Analysis
CulturalVQA comprises 2,328 unique images. In Fig. 2 we show representative samples. We choose images to ensure significant cultural representation across $1$ different countries. The distribution of unique image count per country is detailed in Fig. 3 (A).
@@
We collect 2,378 questions in total. In Fig. 3 (A), we present the number of unique questions per country. The questions have an average length of 10.98 words (see Fig. 3 (B) for country-wise breakdown). Most frequent question types include ‘What’(51.3%), ‘Which’(11.2%), ‘In’ (5.6%), and ‘Why’ (3.4%) questions. For example, ‘What’ questions often relate to identifying cultural entities like saree or Dirndl (traditional Indian and German dresses, respectively) in the clothing category, or festivals like Spring Festival (celebrated in China) among rituals. ‘Where’ questions inquire about locations significant to specific foods, such as the origins of Quebec chicken. Finally, we analyze whether the collected questions contain stereotypes and found that they are largely absent (see § C).
@@
CulturalVQA consists of 7,206 manually curated answers in total. [^6] We collect 1-5 answers per question, depending on the availability of annotators. The average answer length is 1.73 words (see Fig. 3 (B) for country wise breakdown). We assess whether answers predominantly feature terms from local languages. To this end, we verified how many answers have corresponding English Wikipedia titles; for 80% of the answers at least one of the answer words is contained in at least one Wikipedia title. Thus our benchmark is still suitable for English VLMs.
@@
According to the pie chart in Fig. 4, food-related questions are most prevalent, accounting for 37.3% of the dataset, followed closely by traditions and rituals, which represent 26.1% and 18% respectively. Thus, roughly 44% of the questions in our dataset probe for cultural understanding of the intangible aspects of culture (rituals and traditions). The word clouds generated from the collected answers in Fig. 4 illustrate the diversity of expressions, such as hamam (Turkey) and meskel (Ethiopia) for rituals and traditions, and feijoada (Brazil), fufu (Nigeria), and vada (India) for food, indicating a geographically diverse culinary and cultural scope. While the clothing category is the least prevalent in the dataset, it shows the highest variety in terms of collected answers. The drink category is notably one of the smallest, both in terms of the size and number of unique answers.
5 Evaluating VLMs on CulturalVQA